OpenStudy (anonymous):

The city is installing fencing around a rectangular space for a new dog-walking park. The length of the rectangle is 2.25 times its width. If the length measures 150 yards, what will be the perimeter of the dog-walking park?

OpenStudy (anonymous):

I got 433.33 Yards but can't figure out how they got that answer

OpenStudy (anonymous):

so 2*150=300 and then what?

OpenStudy (anonymous):

sorry, its not correct

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

we have: width= length/5= 150/2.25=...?

OpenStudy (anonymous):

so 150 divided by 2.25 and then what?

OpenStudy (michele_laino):

the requested perimeter is: \[\large L + W + L + W = 150 + W + 150 + W = ...?\] where: \[\large W = \frac{L}{{2.25}} = \frac{{150}}{{2.25}} = ...?\]

OpenStudy (anonymous):

150

OpenStudy (michele_laino):

we have: \[\large W = \frac{L}{{2.25}} = \frac{{150}}{{2.25}} = 66.667\] so: perimeter is: \[\large 150 + W + 150 + W = 300 + 66.667 + 66.667 = ...?\]

OpenStudy (anonymous):

thankyou!!

OpenStudy (michele_laino):

thank you! :)